# Cosmology on the Generalized Proca Theory

### v1

@@ Abstract
Abstract
This thesis employs the dynamical systems approach to investigate two distinct cosmological models: one based on a dark energy scenario in an anisotropic Bianchi-I background, and the other exploring the viability of dark energy or inflation within the Generalized SU(2) Proca theory in an isotropic background. Both models address key challenges in understanding the expansion history of the Universe. In the first part, we develop a numerical framework to analyze anisotropic dark energy scenarios involving the interaction between a scalar tachyon field and a vector field in a Bianchi-I background. This approach circumvents the limitations of analytical fixed points, enabling the identification of parameter regions where anisotropic accelerated attractor solutions exist, and providing a versatile tool for investigating dark energy models.
The second part focuses on the Generalized SU(2) Proca (GSU2P) theory, an extension of vector-tensor theories recognized for its versatility in describing phenomena such as inflation, late-time acceleration, and compact objects. Using the dynamical systems approach, we evaluate its potential to drive both primordial inflation and late-time cosmic acceleration within a flat FLRW background. Our analysis reveals significant limitations of the GSU2P model, particularly the absence of stable attractors and smooth transitions between cosmological epochs, caused by indeterminacies in the expansion rate of the Universe, which either diverges to infinity or takes on complex values. While the method identifies theoretical regions of interest, it ultimately rules out the model as a complete explanation for the expansion history of the Universe and as a comprehensive description of inflation with a graceful exit to the radiation-dominated era.
Overall, this work demonstrates the utility of dynamical systems techniques in probing the viability of alternative cosmological models and emphasizes the importance of refining theoretical frameworks to reconcile with observational data. Despite the challenges uncovered, the insights gained here contribute to advancing our understanding of the dynamical evolution of the Universe.
@@
Dark energy, dynamical system approach, accelerated expansion.
@@ Acknowledgements.
I would like to express my deepest gratitude to my family — my father Santiago Garcia Restrepo, my mother Maria Teresa Serna Diaz, and my sister Maria Camila Garcia Serna — for their unwavering support, patience, and encouragement throughout these intense years of academic training. Their readiness to assist me and provide everything I needed has been invaluable. I would like to extend a special thank you to J. Bayron Orjuela-Quintana for his invaluable support and insightful discussions throughout the research process. His contributions were instrumental to the development of this project. I also wish to express my sincere gratitude to my professors, especially Prof. César Alonso Valenzuela, for his unwavering academic and professional guidance, and to Prof. Yeinzon Rodriguez for his continuous support and encouragement. Finally, I am deeply grateful to Universidad del Valle and the Department of Physics for their financial support over the past two years. Their assistance made it possible for me to pursue my post-graduate studies.
@@
Modern cosmology has provided us with the tools to determine the age of the Universe, understand the formation of its structures, discern its energetic composition, and track its evolution over time [1, 2, 3]. These advancements are built upon the cosmological principle [^1] The cosmological principle asserts that the Universe is statistically homogeneous and isotropic at large scales [1]. and the $\Lambda$CDM model, which remains the standard framework in cosmology. In this model, the cosmological constant $\Lambda$ accounts for the accelerated expansion of the Universe [4], while cold dark matter (CDM) is responsible for the formation of cosmic structures [1]. However, despite its successes, recent cosmological data have revealed discrepancies that challenge the adequacy of the $\Lambda$CDM model, such as the Hubble tension [5, 6, 1, 7], the $S_{8}$ tension [8, 9, 10, 11], anomalies in the cosmic microwave background (CMB) [12, 13], and the cosmic dipole misalignment [14, 15, 16], among others [17, 18, 19]. These findings suggest that it is time to explore alternatives to the $\Lambda$CDM framework.
In response to these tensions, several theoretical models have emerged, proposing modifications to the standard cosmological model. These alternatives generally fall into two categories: dynamical dark energy models, including quintessence, k-essence, and tachyonic fields [20, 21, 22, 23, 24, 25], and covariant modifications of gravity, such as $f(R)$ gravity, theories with non-minimal couplings to gravity, and covariant Galileons [26, 27, 28, 29].
Two general classes of models encompass many of these alternatives: one based on scalar fields, known as Horndeski theory [30, 31], and the other based on vector fields, referred to as generalized Proca theory [32, 33, 34]. The Proca theory, in particular, has been shown to be highly adaptable to a wide range of cosmological scenarios. It has been employed as a source of inflation during the early Universe [35, 36] and as a candidate for dark energy in the late Universe [37]. Moreover, Proca theory is inspired by the behavior of fundamental fields in particle physics, allowing for modifications in the internal symmetries of the field [38] or in its interactions with the dark sector [39, 40].
A powerful approach to analyze the dynamics of these models is the use of dynamical system techniques [20, 41, 42], which transform the second-order differential equations of motion into a set of autonomous first-order differential equations. The stationary solutions to these equations reveal the behavior of the system in phase space. This method has been widely applied in cosmology [43, 23, 24, 44, 45, 37, 46].
In this thesis, we will show how to implement the numerical systems approach to an anisotropic dark energy model based on tachyon fields, identifying the proper conditions to reproduce the evolution history of the expanding Universe. Furthermore, we apply the dynamical systems approach to investigate the generalized SU(2) Proca (GSU2P) theory in a flat, homogeneous, and isotropic background. Our goal in this part is to explore the conditions under which the GSU2P theory can drive cosmic acceleration, either during the early inflationary phase or the late-time accelerated expansion. We identify one fixed point that represents de Sitter expansion, which may be a stable or transient state in the evolution of the Universe, along with other "pseudo-stationary" states at distinct scales in the two-dimensional phase space, corresponding to accelerated expansion and radiation-like behavior.
1.1 Background of Standard Cosmology Model $\Lambda$CDM - General Relativity
The formulation of the current standard cosmological model stems from profound questions that have fascinated humanity for centuries, often beginning in childhood: “Why are we here, and how did we come to be here?” Over the past century, these age-old curiosities have transformed into more precise scientific inquiries: “What are the fundamental components of the Universe?”, “Why is the Universe so spatially flat?”, and “How did galaxies form?” [47]. These questions form the core of modern cosmology, shaping the development of the standard model, which is deeply rooted in the principles of general relativity and field theory. Central to this model are Einstein’s field equations, which describe the intricate relationship between the geometry of space-time and the distribution of energy and matter within it,
$R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}+\Lambda g_{\mu\nu}=\frac{1}{m_{\text{P}}^{2}}T_{\mu\nu},$ (1.1)
@@ 1.1 Background of Standard Cosmology Model $\Lambda$CDM - General Relativity
where $R_{\mu\nu}$ denotes the Ricci tensor, $R$ is the scalar of curvature, $g_{\mu\nu}$ represents the metric tensor, $\Lambda$ is the cosmological constant, $m_{\text{P}}$ is the reduced Planck mass, and $T_{\mu\nu}$ is the energy-momentum tensor. Throughout this thesis, natural units $c=\hbar=k_{B}=1$ are employed.
1.1.1 Introduction to the Large-Scale Description of the Universe
Scale Factor
The Universe today is expanding, that is, the average distance between galaxies is much greater today than it was in the past, and this expansion is also accelerating [48, 4]. For this reason, it is better to have cosmological descriptions with respect to background or commoving coordinates such that the physical coordinate is:
$\vec{r}=a(t)\vec{x},$ (1.2)
where $a(t)$ accounts for the expansion of the Universe and is known as the scale factor, this can be schematized as seen in Figure 1.1.
Figure 1.1: Relationship between the co-moving coordinates, blue dots, and physical coordinates, the side of the square at different times.
In this framework, comoving distances remain constant, while physical distances change over time due to the expansion. Similarly, the velocities of particles are described by:
$\dot{\vec{r}}=\frac{\dot{a}}{a}\vec{r}+a\dot{\vec{x}}=H\vec{r}+\vec{v}_{p},$ (1.3)
where $H$ is known as the Hubble parameter and accounts for the rate of expansion of the Universe, $\vec{v}_{p}$ is the peculiar velocity, or the velocity in the background, but on average at cosmological scales this velocity is zero, as evidenced in Figure 1.2.
Figure 1.2: Hubble diagram for measuring the recession velocity for different objects as a function of distance, the slope accounts for the rate of expansion or Hubble parameter today, $H_{0}$, taken from [47].
Thus, we can approximate the relationship as
@@ 1.1.1 Introduction to the Large-Scale Description of the Universe
where this equation is known as the Hubble law and $H_{0}$ denotes the Hubble parameter today, also called the Hubble constant[49].
Redshift
One of the primary methods for studying cosmological phenomena, measuring distances, and determining the composition of celestial bodies relies on the detection and analysis of electromagnetic radiation[47]. The frequency of this radiation, however, is altered not only by the relative motion between the source and the observer but also significantly by the expansion of the Universe itself. As the Universe expands, the wavelengths of emitted radiation stretch proportionally to the scale factor $a(t)$. Thus, if an object emits radiation with an initial wavelength $\Delta\lambda_{0}$, the observed wavelength on Earth becomes $\Delta\lambda_{f}=a(t)\lambda_{0}$. This stretching of wavelengths, known as redshift, provides a direct measure of the expansion of the universe, with the redshift factor defined as
$1+z_{r}\equiv\frac{\Delta\lambda_{f}}{\Delta\lambda_{0}}=\frac{a(t_{0})}{a(t)},$ (1.5)
where $z_{r}$ is the redshift, and it is conventional to set the scale factor at the present time as $a(t_{0})=1$. It is important to note that if $z_{r}>0$, the detected wavelength is longer than the emitted one (indicating the object is going out), hence the term redshift. Conversely, if $z_{r}<0$, the detected wavelength is shorter than the emitted one (indicating the object is coming in), which is referred to as blueshift. For very distant objects, however, the peculiar velocity is nearly negligible, as illustrated in Figure 1.2, allowing for an approximation where redshift is directly related to the time elapsed since the observed event occurred.
Cosmological Principle and Geometry
In formulating a theory of the Universe, it is crucial to incorporate observational symmetries, leading to the postulate that the Universe possesses the maximum possible number of symmetries, or equivalently, behaves as a de-Sitter Universe. The primary justification for this postulate comes from observations of the Cosmic Microwave Background (CMB), which reveal that the Universe is statistically homogeneous and isotropic on large scales, with anisotropies of the order of $\Delta T\sim 10^{-5}\,\text{K}$, a value not sufficiently precise to challenge this assumption as it falls within the observational error range [1]. Therefore, if the Universe is homogeneous and isotropic, the most general metric that can describe this geometry is the Friedman-Lemaître-Robertson-Walker (FLRW) metric, which must reduce to the Minkowski metric of special relativity in the flat limit, $g_{\mu\nu}\sim\text{diag}(-1,1,1,1)$.
To explicitly demonstrate homogeneity and isotropy, the spatial part of the metric can be considered as the implicit metric of a 4D hypersphere, with radius $R$ or $iR$:
@@
$dw$ $=-\frac{r\,dr}{\sqrt{\sigma R^{2}-r^{2}}},$ (1.8)
where $\sigma=\{-1,0,1\}$ and $r^{2}=x^{2}+y^{2}+z^{2}$. Given that distances are measured in a 4D Euclidean space with a constraint, the spatial line element is
$d\Sigma^{2}$ $=dw^{2}+dx^{2}+dy^{2}+dz^{2},$ (1.9)
$d\Sigma^{2}$ $=\frac{r^{2}\,dr^{2}}{\sigma R^{2}-r^{2}}+dr^{2}+r^{2}d\Omega,$ (1.10)
$d\Sigma^{2}$ $=\frac{\sigma R^{2}\,dr^{2}}{\sigma R^{2}-r^{2}}+r^{2}d\Omega,$ (1.11)
which leads to the FLRW line element, incorporating the temporal component, $ds^{2}=-dt^{2}+a^{2}(t)d\Sigma^{2}$:
$ds^{2}=-dt^{2}+a^{2}(t)\left(\frac{dr^{2}}{1-kr^{2}}+r^{2}d\Omega^{2}\right),$ (1.12)
where $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta\,d\phi^{2}$ is the differential solid angle in spherical coordinates, $t$ represents cosmic time, $r$ is the comoving distance, $a(t)$ is the scale factor, and $k=\sigma/R^{2}$ is the spatial curvature. Depending on the value of $k$, the Universe can be classified as spatially open ($k<0$), spatially closed ($k>0$), or spatially flat ($k=0$), as illustrated in Figure 1.3.
Figure 1.3: Diagram illustrating the behavior of spatial geometry based on the curvature parameter $k$. The left graph shows a closed, spherical geometry for $k>0$; the center graph represents a flat, Euclidean geometry for $k=0$; and the right graph depicts an open, hyperbolic geometry for $k<0$.
Another Interesting Geometry - Bianchi I
An alternative to the isotropic cosmological background is the anisotropic Bianchi I metric, which describes a universe where the expansion rates along different axes can differ. This metric is particularly useful in scenarios where the cosmological principle may not hold exactly at all scales or if we consider anisotropic anomalies enough to not guarantee the cosmological principle[14, 15, 16, 50]. Unlike the FLRW metric, the Bianchi I metric has a more general form, where the spatial components exhibit expansion with rotational symmetry only in the $y$-$z$ plane, as illustrate in the Figure 1.4. In this context, the Bianchi I metric can be written as:
$\text{d}s^{2}=-\text{d}t^{2}+a^{2}(t)\left[e^{-4\sigma(t)}\text{d}x^{2}+e^{2\sigma(t)}\left(\text{d}y^{2}+\text{d}z^{2}\right)\right],$ (1.13)
where $a(t)$ represents the average scale factor, and $\sigma(t)$ denotes the geometric shear, both of which are functions of the cosmic time $t$. This metric allows for anisotropic behavior during the evolution of the Universe. However, even though we include an anisotropic shear $\sigma(t)$, it must remain small enough to be consistent with observations. This ensures that the anisotropic effects are present but not strong enough to fundamentally challenge the cosmological principle [51].
Figure 1.4: Diagram illustrating the evolution of a comoving perfect sphere under the Bianchi-I metric, Eq. (1.13). The red sphere represents the initial condition. Over time, the sphere deforms into an ellipse, as shown by the blue ellipse, and further evolves into a more elongated and larger shape, represented by the black ellipse.
1.1.2 Energy-Momentum Tensor
The most general energy-momentum tensor for any fluid is

### v2

@@ Abstract
In this thesis, we employ the dynamical system approach to investigate two cosmological models: the first one based on a dark energy scenario in an anisotropic Bianchi-I background, and the other exploring the viability of dark energy or inflation in an isotropic background. Both models address key challenges in understanding the expansion history of the Universe. In the first part, the complexity introduced by shear prevents the analytical identification of certain fixed points in the model. To address this, we develop a numerical framework to study anisotropic dark energy scenarios involving the interaction between a scalar tachyon field and a vector field in a Bianchi-I background. This approach overcomes the limitations of analytical fixed-point methods, allowing us to identify parameter regions that support anisotropic accelerated attractor solutions and providing a versatile tool for exploring dark energy models.
The second part focuses on the Generalized SU(2) Proca theory, an extension of vector-tensor theories which has recently garnered attention for its potential to describe key phases of cosmic evolution, including primordial inflation and late-time accelerated expansion. However, its full cosmological implications remain unexplored. In this part, we perform a comprehensive analysis of the dynamical properties of the GSU2P theory in a flat, homogeneous, and isotropic spacetime, through a dynamical-system approach. Our analysis reveals the presence of three pairs of fixed points, one of them corresponding to de-Sitter expansion which may represent either a stable or unstable phase in the evolution of the universe. These points, nonetheless, give rise to an indeterminate or infinite Hubble parameter, which renders them cosmologically unviable. Additionally, we find two key pseudostationary states: the attractor lines, along which the system exhibits constant-roll dynamics, and the central zone, characterized by oscillatory radiation-like behaviour of the field. The dynamics within the central zone could represent a graceful exit from the primordial inflationary phase to a radiation dominated phase, or a state of the dark energy component prior to the late-time cosmic acceleration. However, within the central zone, the dynamics of the vector field leads to recurrent instances of a nonphysical expansion rate. The absence of a limit cycle in the central zone further exacerbates the issue, as the system may follow unbounded phase-space trajectories, and the expansion rate becomes complex once it escapes the region. Collectively, these challenges undermine the viability of the GSU2P theory as a cosmological model for cosmic acceleration.
Overall, this work employs dynamical systems techniques to assess the viability of alternative cosmological models. Our numerical framework circumvents analytical limitations in anisotropic dark energy scenarios, identifying parameter regions that allow for accelerated expansion. For the Generalized SU(2) Proca theory, we demonstrate that despite the presence of pseudostationary structures suggesting possible cosmological transitions, the model ultimately fails due to recurrent nonphysical expansion rates and unstable trajectories.
@@
Dark energy, dynamical system, accelerated expansion, inflation.
@@ Acknowledgements.
I would like to express my deepest gratitude to my family — my father Santiago García Restrepo, my mother Maria Teresa Serna Díaz, and my sister Maria Camila García Serna — for their unwavering support, patience, and encouragement throughout these intense years of academic training. Their readiness to assist me and provide everything I needed has been invaluable. I would like to extend a special thank you to J. Bayron Orjuela-Quintana for his invaluable support and insightful discussions throughout the research process. His contributions were instrumental to the development of this project. I also wish to express my sincere gratitude to my professors, especially Prof. César Alonso Valenzuela, for his unwavering academic and professional guidance, and to Prof. Yeinzon Rodriguez for his continuous support and encouragement. Finally, I am deeply grateful to Universidad del Valle and the Department of Physics for their financial support over the past two years. Their assistance made it possible for me to pursue my post-graduate studies.
@@
Einstein’s General Relativity (GR) revolutionized our comprehension of the Universe, replacing centuries of philosophical speculation with a mathematically rigorous framework for describing the dynamics of spacetime. This groundbreaking theory redefined gravity not as a force acting at a distance, but as the geometric manifestation of spacetime curvature induced by the presence of matter and energy. With this profound conceptual shift, General Relativity laid the cornerstone for modern cosmology. The Einstein field equations encode the fundamental interplay between the geometry of spacetime and the distribution of matter-energy within the Universe, expressed as:
$R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}+\Lambda g_{\mu\nu}=8\pi GT_{\mu\nu},$ (1.1)
@@
where $R_{\mu\nu}$ represents the Ricci tensor, $R$ is the Ricci scalar, $g_{\mu\nu}$ denotes the metric tensor, $T_{\mu\nu}$ is the energy-momentum tensor, $\Lambda$ is the cosmological constant, and $m_{\text{P}}$ is the reduced Planck mass. Throughout this work, we adopt natural units where $c=\hbar=k_{B}=1$.
For the first time in history, the most fundamental cosmological inquiries could be addressed through the rigorous lens of physical law: What are the origins of the light elements? Why does the Universe exhibit such remarkable homogeneity and isotropy on large scales? What mechanisms were responsible for generating the primordial density fluctuations that subsequently evolved into galaxies and the intricate cosmic web? These profound questions, once confined to the realm of metaphysical speculation, have found compelling, albeit partial, answers through the synergistic interplay between theoretical physics and increasingly precise astronomical observations.
The modern cosmological framework is firmly rooted in the paradigm of an expanding Universe, a picture robustly supported by a wealth of empirical evidence [1, 2, 3]. This model traces the history of the cosmos back to an exceedingly hot and dense primordial state. The subsequent evolution of the Universe is described with remarkable accuracy by the synthesis of General Relativity and the Standard Model of particle physics, augmented by three crucial components: cold dark matter, invoked to explain the formation of large-scale structures; dark energy, introduced to account for the observed accelerated expansion at late times; and an early inflationary epoch, postulated to resolve the horizon and flatness problems and to provide a mechanism for the generation of primordial perturbations.
1.1 Measuring Distances in the Universe
Figure 1.1: Schematic representation of cosmic expansion. Two points with fixed comoving coordinates maintain a constant comoving separation, while their physical distance grows proportionally to $a(t)$, illustrating the metric expansion of spacetime.
1.1.1 Scale Factor
The current Universe is characterized by an accelerated expansion, a phenomenon robustly confirmed through observations of Type Ia supernovae [1, 2, 3]. This expansion implies that the average physical separation between gravitationally unbound structures, such as galaxies or galaxy clusters, increases over cosmic time. To describe this large-scale dynamic without being tied to the stretching fabric of spacetime itself, cosmologists employ comoving coordinates, $\mathbf{x}$. This system factors out the universal expansion by introducing a dimensionless, time-dependent scale factor, $a(t)$. Within this framework, the physical position $\mathbf{r}(t)$ of an object is related to its fixed comoving coordinate $\mathbf{x}$ by:
$\mathbf{r}(t)=a(t)\mathbf{x}.$ (1.2)
By convention, the scale factor is normalized such that $a(t_{0})=1$ at the present cosmic time $t_{0}$. Figure 1.1 provides a schematic illustration: as the Universe evolves, commoving coordinates (grid points) remain fixed, while the physical distances between them scale proportionally with $a(t)$, reflecting the metric expansion.
The kinematics of cosmic expansion can obtain examining the time derivative of the physical position $\mathbf{r}(t)$. Defining the Hubble parameter $H(t)\equiv\dot{a}(t)/a(t)$ as the rate of expansion, the total velocity of a particle relative to an observer at the origin is given by:
$\dot{\mathbf{r}}=\dot{a}\mathbf{x}+a\dot{x}=H(t)\mathbf{r}+\mathbf{v}_{p}.$ (1.3)
This expression decomposes the total velocity into two components:
  1. 1.
    The Hubble flow, $H(t)\mathbf{r}$, which represents the velocity due to the overall expansion of space.
  2. 2.
    The peculiar velocity, $\mathbf{v}_{p}\equiv a(t)\dot{\mathbf{x}}$, which describes the motion of the object relative to the comoving grid, typically driven by local gravitational interactions.
Figure 1.2: A modern Hubble diagram illustrating the linear relationship between the recession velocities (often inferred from redshift) and distances of galaxies. The slope of the best-fit line corresponds to the Hubble constant, $H_{0}$. Data shown are representative of modern measurements, such as those from [3], reinforcing this fundamental cosmological principle.
On sufficiently large scales, the effects of local structures average out, making peculiar velocities statistically negligible ($\langle\mathbf{v}_{p}\rangle\approx 0$). In this regime, the motions of distant objects are dominated by the homogeneous Hubble flow. This theoretical picture directly leads to the empirically observed Hubble-Lemaître law:
@@ 1.1.1 Scale Factor
where $v$ is the recession velocity of a distant galaxy, $r$ is its proper distance, and $H_{0}=H(t_{0})$ is the Hubble constant, the value of the Hubble parameter today [4]. The linearity of this relationship is confirmed by redshift-distance measurements like those shown in Figure 1.2.
1.1.2 Redshift as a Consequence of Expansion
While the scale factor $a(t)$ provides the theoretical framework for cosmic expansion, its most direct and crucial observable consequence is the effect it has on electromagnetic radiation traveling across cosmological distances. As photons propagate through the expanding Universe from a distant source to an observer, the stretching of spacetime itself causes their wavelengths to increase. This phenomenon is known as cosmological redshift [5].
Consider light emitted by a source at cosmic time $t_{\text{em}}$ with a wavelength $\lambda_{\text{em}}$. As this light travels towards us, the Universe continues to expand. By the time it is detected at the present time $t_{0}$, its wavelength will have been stretched to $\lambda_{\text{obs}}$. The amount of stretching is directly proportional to the change in the scale factor during the light’s journey:
$\frac{\lambda_{\text{obs}}}{\lambda_{\text{em}}}=\frac{a(t_{0})}{a(t_{\text{em}})}.$ (1.5)
Cosmological redshift, denoted by $z$, is defined as the fractional change in wavelength:
$z\equiv\frac{\lambda_{\text{obs}}-\lambda_{\text{em}}}{\lambda_{\text{em}}}=\frac{\lambda_{\text{obs}}}{\lambda_{\text{em}}}-1.$ (1.6)
Combining these equations and using the convention $a(t_{0})=1$, we arrive at the fundamental relationship between redshift and the scale factor at the time of emission:
$1+z=\frac{a(t_{0})}{a(t_{\text{em}})}=\frac{1}{a(t_{\text{em}})}.$ (1.7)
Therefore, measuring the redshift $z$ of a distant object provides a direct measurement of the scale factor of the Universe when the light was emitted ($a(t_{\text{em}})=1/(1+z)$). Since $\lambda_{\text{obs}}>\lambda_{\text{em}}$ for an expanding universe ($a(t_{0})>a(t_{\text{em}})$), $z$ is positive, indicating a shift towards longer (redder) wavelengths—hence the term “redshift”.
It is important to distinguish this cosmological redshift, caused by the expansion of space itself, from the Doppler redshift caused by the peculiar velocity of the source relative to the observer. While both effects can alter observed wavelengths, for distant objects ($z\gtrsim 0.1$), the cosmological component typically dominates and serves as the primary indicator of distance (via the Hubble-Lemaître law) and cosmic epoch [5, 6]. Redshift measurements are thus fundamental not only for mapping the large-scale structure of the Universe but also for probing its expansion history and composition.
1.1.3 The Geometric Framework: Isotropy and the FLRW Metric
The scale factor $a(t)$ and cosmological redshift $z$, discussed previously, describe the dynamics of cosmic expansion. However, a complete cosmological model requires specifying the geometric stage upon which this expansion unfolds. The standard model of cosmology is built upon the Cosmological Principle, a foundational assumption postulating that the Universe is statistically homogeneous and isotropic on sufficiently large scales. This principle is strongly supported by observations, particularly of the Cosmic Microwave Background (CMB), which reveals a remarkably uniform temperature across the sky, with anisotropies only at the level of $\Delta T/T\sim 10^{-5}$ [7, 8, 9, 10].
Symmetries, like those encapsulated in the Cosmological Principle, are useful in physics; they simplify descriptions and often reveal fundamental properties. For a spacetime exhibiting spatial homogeneity and isotropy, the most general line element describing its geometry is the Friedman-Lemaître-Robertson-Walker (FLRW) metric. In the limit of no gravity or expansion ($a(t)=1,k=0$), it reduces to the Minkowski metric $g_{\mu\nu}=\text{diag}(-1,1,1,1)$ of special relativity.
Figure 1.3: Diagram illustrating the three possible spatial geometries in the FLRW model, determined by the curvature parameter $k$. Left ($k>0$): A closed, spherical geometry where parallel lines eventually converge. Center ($k=0$): A flat, Euclidean geometry where parallel lines remain parallel. Right ($k<0$): An open, hyperbolic geometry where parallel lines diverge.
The spatial part of the FLRW metric can be derived by considering a 3-dimensional hyperspace embedded in a 4-dimensional Euclidean or pseudo-Euclidean space. Depending on whether the embedding space has a positive, zero, or negative curvature signature ($\sigma=+1,0,-1$, respectively), the geometry of the 3D slice changes. Let $R$ be the characteristic radius of curvature (infinite for $\sigma=0$). A point in the embedding space satisfies:
@@ 1.1.3 The Geometric Framework: Isotropy and the FLRW Metric
$dw$ $=-\frac{r\,dr}{w}=-\frac{r\,dr}{\sqrt{\sigma R^{2}-r^{2}}}.$ (1.10)
The line element $d\Sigma^{2}$ in the 3D spatial slice is inherited from the 4D embedding space’s metric $d\Sigma^{2}=dx^{2}+dy^{2}+dz^{2}+dw^{2}$. Using spherical coordinates for $x,y,z$ ($dr^{2}+r^{2}d\Omega^{2}=dx^{2}+dy^{2}+dz^{2}$), we get:
$d\Sigma^{2}$ $=dr^{2}+r^{2}d\Omega^{2}+dw^{2}$ (1.11)
$=dr^{2}+r^{2}d\Omega^{2}+\frac{r^{2}\,dr^{2}}{\sigma R^{2}-r^{2}}$ (1.12)
$=\frac{\sigma R^{2}}{\sigma R^{2}-r^{2}}dr^{2}+r^{2}d\Omega^{2}.$ (1.13)
Defining the curvature parameter $k=\sigma/R^{2}$ where $k$ can take values $\{k<0,k=0,k>0\}$, we can write:
$d\Sigma^{2}=\frac{dr^{2}}{1-kr^{2}}+r^{2}d\Omega^{2}.$ (1.14)
Incorporating the time dimension and the expansion via the scale factor $a(t)$, we arrive at the full FLRW line element:
$ds^{2}=-dt^{2}+a^{2}(t)\left[\frac{dr^{2}}{1-kr^{2}}+r^{2}\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right)\right],$ (1.15)
where $t$ is the cosmic time, $r,\theta,\phi$ are comoving spherical coordinates and $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta\,d\phi^{2}$. The parameter $k$ dictates the global geometry of the spatial sections:
  • $k>0$: Positive curvature, corresponding to a closed, finite Universe (like a 3-sphere).
  • $k=0$: Zero curvature, corresponding to a flat, infinite Euclidean Universe.
  • $k<0$: Negative curvature, corresponding to an open, infinite hyperbolic Universe.
These possibilities are illustrated schematically in Figure 1.3. Current observations, primarily from the CMB, strongly favour a spatially flat Universe ($k\approx 0$[7, 10].
1.1.4 The Cosmic Fluid: Energy-Momentum Tensor and Equations of State
Having established the standard geometric framework for cosmology – the FLRW metric (1.15) based on the Cosmological Principle – we now turn to the source term in Einstein’s field equations: the energy and matter content of the Universe. This content dictates the expansion dynamics $a(t)$ within the given geometry. The distribution of energy and momentum is described by the energy-momentum tensor, $T_{\mu\nu}$.
In its most general form for any continuous medium or field configuration, $T_{\mu\nu}$ is a symmetric rank-2 tensor whose components represent densities and fluxes of energy and momentum. In a local inertial frame, its components have direct physical interpretations: